OpenStudy (anonymous):

maggies mold has a right cylindrical shape. the mold is 8 inches high, and the bottom of the mold has a diameter of 5 inches. how many cubic inches are in maggies mold?

OpenStudy (saifoo.khan):

Volume = base area * height.

OpenStudy (anonymous):

doesn't it need to include pi?

OpenStudy (saifoo.khan):

for base area yes, you will have to add.

OpenStudy (anonymous):

ok so the formula is V=b*a*h?

OpenStudy (saifoo.khan):

(base area) * h \[\pi r^2 * height\]

OpenStudy (anonymous):

would the answer be 70.50 cubic inches?

OpenStudy (saifoo.khan):

it will be, pi* (2.5)^2 * 8 50 pi

OpenStudy (anonymous):

will the formula be V=pi*H? where would the diameter go?

OpenStudy (anonymous):

thank you i get the problem now :)

OpenStudy (saifoo.khan):

radius is half of diameter.

OpenStudy (anonymous):

yes. i know. i get how you got 50 pi

OpenStudy (saifoo.khan):

cool. :D

OpenStudy (saifoo.khan):

please hit good answer.